OpenStudy (anonymous):

A screw has a mechanical advantage of 8. How much force must you apply to the screw so that it has an output force of 750 N? A. 550 N B. 94 N C. 55 N D. 6000 N

OpenStudy (jfraser):

\[MA = \frac{force \space output}{force \space input}\]you know the advantage and the neded output force, solve for the input force

OpenStudy (anonymous):

I worked that out and it was wrong, the answer was B.

OpenStudy (jfraser):

then you worked it out wrong. If you rearrange to solve for input, then\[force input = \frac{force output}{MA} = \frac{750N}{8} = 93.75N \approx 94N\]